sam (.sam.):

A. when multiplying or dividing exponents with the same base (which is 2 in this case) all you have to do is add or subtract the two exponents together on top and keep the base. So for this problem, we are multiplying 2^20 by 2^3. add 20 to 3 and get 2^23 as your final answer for amount of bits in a megabyte. B. 2^30 divided by 2^20 = 2^10 for megabytes in a gb. 2^3 times 2^30 = 2^33

OpenStudy (anonymous):

megabyte (mb) is 2^20 bytes -> Bit = 2^20 * 2^3 = 2^23 bits Gigabyte = 10^30 bytes -> Meg = 2^30/ 2^20 = 2^10 Mega bytes -> Bit = 2^30 * 2^3 = 2^33 bits